# Contrastive Learning and Adversarial Disentanglement for Task-Oriented Semantic Communications

### v2

@@ Abstract
Task-oriented semantic communication systems have emerged as a promising approach to achieving efficient and intelligent data transmission, where only information relevant to a specific task is communicated. However, existing methods struggle to fully disentangle task-relevant and task-irrelevant information, leading to privacy concerns and subpar performance. To address this, we propose an information-bottleneck method, named CLAD (contrastive learning and adversarial disentanglement). CLAD utilizes contrastive learning to effectively capture task-relevant features while employing adversarial disentanglement to discard task-irrelevant information. Additionally, due to the lack of reliable and reproducible methods to gain insight into the informativeness and minimality of the encoded feature vectors, we introduce a new technique to compute the information retention index (IRI), a comparative metric used as a proxy for the mutual information between the encoded features and the input, reflecting the minimality of the encoded features. The IRI quantifies the minimality and informativeness of the encoded feature vectors across different task-oriented communication techniques. Our extensive experiments demonstrate that CLAD outperforms state-of-the-art baselines in terms of semantic extraction, task performance, privacy preservation, and IRI. CLAD achieves a predictive performance improvement of around 2.5-3%, along with a 77-90% reduction in IRI and a 57-76% decrease in adversarial attribute inference attack accuracy.
@@
Contrastive learning, disentangled representation learning, information-bottleneck, semantic communications, task-oriented communications.
@@ I Introduction
In conventional communication systems, the primary objective has been to ensure reliable transmission of data, focusing on delivering bit sequences across noisy channels without considering the meaning, context, or purpose of the data being transmitted. Shannon’s mathematical theory of communication focuses on optimizing metrics such as data rate, error rate, and bandwidth efficiency, whilst being agnostic to the ultimate purpose and relevance of the transmitted information [1]. This approach has been widely successful and effective for general communication needs thus far. However, next-generation communication systems, beginning with 6G, require more intelligent and task-aware communication methods to enable emerging applications [2, 3], such as computer vision [4], autonomous driving [5], extended reality (XR) [6], and generative artificial intelligence (AI) [7].
As we move towards these advanced systems, there is a growing recognition that communication should not merely be about transmitting raw data, but about understanding and leveraging the underlying meaning and purpose of the data. This shift towards task-oriented semantic communications represents a fundamental change in the design of communication networks [8, 9]. Instead of focusing solely on the accurate and efficient transmission of bits, these new approaches aim to ensure that the information most relevant to the specific task or decision-making process is prioritized and delivered with minimal delay. For example, in a smart city environment [10], rather than transmitting all sensor data from traffic cameras, task-oriented communications focuses on sending only the information necessary to identify and respond to potential hazards or optimize traffic flow in real time.
With the growing success and popularity of deep learning (DL) in various wireless communication applications [11, 12], many emerging task-oriented communication systems have adopted DL approaches to encode task-relevant information to improve task performance and efficiency of the communication system[13, 14, 15]. Nevertheless, most proposed schemes do not focus quantifying or benchmarking the amount of information that the encoded features retain about the input, primarily due to the computational difficulty of estimating mutual information and the lack of a unified methodology that provides fair and reproducible results. This omission is critical, as understanding the mutual information is key to developing systems with improved bandwidth and latency. Moreover, there has been little emphasis on explicitly discarding task-irrelevant information, often leading to a trade-off between task performance and information compression.
@@
To address the aforementioned challenges, we develop a task-oriented communication system based on contrastive learning [17] and disentangled representation learning [18] and we devise a new metric to compute comparative values for a proxy of mutual information between the encoded features and the inputs across different systems, rather than computing the exact mutual information. More specifically, our major contributions are as follows:
  • We derive a lower bound for the mutual information between the encoded features and the target using contrastive learning principles. We show that the contrastive learning based lower bound improves task accuracy and performance compared to the traditional cross-entropy based mutual information approximations;
  • We propose a systematic training methodology based on an innovative loss function, designed to extract task-irrelevant information through reconstruction losses while disentangling it from task-relevant features using adversarial methods. This enables the system to prioritize the transmission of task-relevant features while minimizing communication overhead and reducing unnecessary information transmission;
  • To address the current limitation of lacking a reliable and unified approach for estimating the mutual information between the encoded features and input data, we introduce a new metric named the information retention index (IRI) which serves as a proxy for the mutual information. This metric compares the informativeness and minimality of the encoded features across various task-oriented communication methods, providing deeper insights into system behavior and enabling a more rigorous comparison of their performance;
  • We evaluate our proposed task-oriented communication system in diverse channel conditions. It is tested against several existing task-oriented communication methods, demonstrating improved task performance, enhanced privacy awareness, and reduced amount of stored information across a wide range of transmission scenarios.
The rest of this paper is organized as follows. In Section II, we provide an overview of related works in task-oriented communications, contrastive learning, and disentangled representation learning. Section III introduces the system model followed by an in-depth discussion of the problem statement and involved challenges. Section V discusses our proposed method (CLAD) and the associated mathematical formulations and algorithms. In Section VI, we outline our experimental setup, and provide a thorough analysis and discussion of the results and their implications. Finally, Section VII concludes the paper with a discussion on the importance of our results as well as future work and possibilities in the realm of task-oriented semantic communications.
@@ II Related Work
DL-based communication systems have shown success in recent years. DeepJSCC (Deep Joint Source-Channel Coding) is a recent advancement in the field of wireless communications that utilizes deep learning to jointly optimize source and channel coding, which are traditionally treated as separate tasks [14]. Unlike conventional methods that rely on separate compression and error-correction codes, DeepJSCC uses neural networks to directly map source data to channel symbols, allowing for an end-to-end optimization of the communication system. DeepJSCC can be trained on a classification task by minimizing the cross-entropy loss, ensuring task-specific performance; however, it does not inherently ensure that only task-relevant features are transmitted.
@@
Wang et al. formulate a privacy-utility trade-off to develop an IB-based privacy-preserving task-oriented communication system against model inversion attacks [20]. This is achieved by striking a balance between the traditional IB-based loss functions similar to the work discussed above and a mean squared error (MSE) based term that aims at maximizing reconstruction distortion. Their results demonstrate improved privacy with minimal impact to task performance.
@@ III-A Notations
Throughout this paper, we use the following notational conventions. Random variables are denoted by uppercase letters, such as $X$, $Y$, and $Z$. Their corresponding realizations (i.e., specific instances) are denoted by lowercase bold letters, such as $\boldsymbol{x}$, $\boldsymbol{y}$, and $\boldsymbol{z}$. The space from which these random variables are drawn is represented by calligraphic letters, such as $\mathcal{X}$, $\mathcal{Y}$ and $\mathcal{Z}$. We denote entropy of a random variable $X$ by $H(X)$. The mutual information between two random variables $X$ and $Y$, is denoted by $I(X;Y)$. We use $I(Z;X|Y)$ to denote the conditional mutual information between $Z$ and $X$ given $Y$. We use the expectation notation $\mathbb{E}[\cdot]$, which refers to the average value of a random variable over a distribution.
@@ III-B System Model
We consider a communication system with a transmitter that includes a feature extractor and a joint source-channel coding (JSCC) encoder. We collectively refer to these components as the task-relevant encoder (TRE). The TRE encodes an input image $\boldsymbol{x}\in\mathcal{X}$ into a lower-dimensional feature vector $\boldsymbol{z}\in\mathcal{Z}$. Encoded vector $\boldsymbol{z}$ is then transmitted to a receiver over a noisy wireless channel. The primary objective is to transmit a minimal yet informative representation ($\boldsymbol{z}$), by discarding task-irrelevant information while ensuring that $\boldsymbol{z}$ contains only the essential information for accurate downstream classification at the receiver.
@@
At the transmitter, the TRE encodes input image $\boldsymbol{x}\in\mathbb{R}^{N}$, where $N$ represents the number of pixels in the image (Height $\times$ Width $\times$ Color Channels). The encoder maps this input into a lower-dimensional feature vector $\boldsymbol{z}\in\mathbb{R}^{d}$, where $d$ is the dimension of the encoded feature vector. The encoding function, denoted by $f_{\theta}:\mathbb{R}^{N}\to\mathbb{R}^{d}$, is parameterized by $\theta$, and the encoding process can be expressed as
@@
Feature vector $\boldsymbol{z}$ is then prepared for transmission over the wireless channel by being mapped to channel input symbols. The role of the TRE is twofold: encoding the input data into feature representations and preparing them as channel symbols suitable for transmission. The encoded feature vector $\boldsymbol{z}\in\mathbb{R}^{d}$ is transmitted over a wireless channel, which is modeled as an additive white Gaussian noise (AWGN) channel. The channel introduces noise and distortion, and the received signal $\hat{\boldsymbol{z}}\in\mathbb{R}^{d}$ at the receiver is expressed as
@@
At the receiver, a classifier, denoted by $q_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{M}$, where $M$ is the number of labels, is parameterized by $\phi$. The classifier maps the received noisy signal $\hat{\boldsymbol{z}}$ to predicted label $\hat{y}\in\mathbb{R}^{M}$ as
@@ IV-D Limitations of Variational Information Bottleneck (VIB)
  • Conflicting objectives: The VIB objective which maximizes $I(\hat{Z};X)$ and minimizes $I(\hat{Z};Y)$ leads to a conflicting objective [31]. If we consider the information diagram [33] presented in Fig. 1, it is evident that the region shaded in blue, namely $I(X;Y;\hat{Z})$ is a subset of $I(\hat{Z};X)$ and $I(\hat{Z};Y)$ and is therefore maximized and minimized simultaneously.
@@ V Proposed Method: CLAD
The core idea of CLAD is to create an end-to-end communication system that explicitly learns to disentangle task-relevant information from task-irrelevant content, enabling both high task accuracy and improved privacy. Rather than simply combining existing techniques, CLAD integrates supervised contrastive learning, adversarial training, and reconstruction-based supervision into a unified, structured framework that aligns with the Minimum Necessary Information (MNI) principle.
@@
To ensure effective optimization, CLAD adopts a three-stage training strategy, where each stage isolates and refines a different component of the system. This design is intentional — it prevents interference between competing objectives, stabilizes training, and encourages modular reuse of the encoders across stages.
CLAD utilizes two key encoders: the TRE, which maps the input into the task-relevant channel codeword $Z_{1}$, and the task-irrelevant encoder (TIE), which maps the input into task-irrelevant channel codeword $Z_{2}$. The task performance is optimized through contrastive learning, which aims to maximize the mutual information $I(\hat{Z_{1}};Y)$, ensuring that $\hat{Z_{1}}$ captures the most informative features for downstream classification. The disentanglement is achieved through reconstruction learning to capture task-irrelevant information in $\hat{Z_{2}}$, and adversarial training is utilized to minimize the mutual information $I(\hat{Z_{1}};\hat{Z_{2})}$, thus promoting independence between the two feature representations. The different components and training stages of CLAD are visualized in Fig. 3. These components are optimized together to ensure both high task accuracy and effective disentanglement of information, corresponding to the following maximization objective:
@@ V-A Contrastive Loss for Task-Relevant Features
where $S_{ij}$ represents the similarity score between the projected representations $\boldsymbol{h}_{i}$ and $\boldsymbol{h}_{j}$, $\tau$ is a temperature scaling factor, and $\mathds{1}_{y_{i}=y_{j}}$ is an indicator function that equals 1 if $y_{i}=y_{j}$ (positive pairs) and 0 otherwise (negative pairs). This loss encourages encoder $f_{\theta}(\cdot)$ to learn class-discriminative features, ensuring that the latent representation $\boldsymbol{z}$ captures the necessary information for the downstream classification task.
@@ V-F Information Retention Index Across Different Methods
While methods such as MINE [36] offer tighter approximations of MI, they require adversarial or variational training, are sensitive to hyperparameters, and are not directly comparable across model architectures. In contrast, IRI allows for consistent, fair benchmarking with minimal assumptions and training overhead. We view IRI as a relative measure, similar to a diagnostic tool, that enables empirical analysis of minimality and informativeness in task-oriented systems. We note that while SSIM is effective for evaluating visual similarity in image-based applications, it is inherently limited to image data. For non-visual modalities such as text or tabular data, alternative reconstruction-based metrics (e.g., BLEU score) can be adopted to extend the IRI framework to broader task domains.

### v3

# Trustworthy and Privacy-Aware Task-Oriented Semantic Communication for 6G-IoT Using Contrastive and Disentangled Representations

@@ Abstract
Task-oriented semantic communication systems have emerged as a promising approach to achieving efficient and intelligent data transmission in next-generation networks, where only information relevant to a specific task is communicated. This is particularly important in 6G-enabled Internet of Things (6G-IoT) scenarios, where bandwidth constraints, latency requirements, and data privacy are critical. However, existing methods struggle to fully disentangle task-relevant and task-irrelevant information, leading to privacy concerns and suboptimal performance. To address this, we propose an information-bottleneck inspired method, named CLAD (contrastive learning and adversarial disentanglement). CLAD utilizes contrastive learning to effectively capture task-relevant features while employing adversarial disentanglement to discard task-irrelevant information. Additionally, due to the absence of reliable and reproducible methods to quantify the minimality of encoded feature vectors, we introduce the Information Retention Index (IRI), a comparative metric used as a proxy for the mutual information between the encoded features and the input. The IRI reflects how minimal and informative the representation is, making it highly relevant for privacy-preserving and bandwidth-efficient 6G-IoT systems. Extensive experiments demonstrate that CLAD outperforms state-of-the-art baselines in terms of semantic extraction, task performance, privacy preservation, and IRI, making it a promising building block for responsible, efficient and trustworthy 6G-IoT services.
@@
Contrastive learning, disentangled representation learning, information-bottleneck, semantic communication, task-oriented communication.
@@ I Introduction
In conventional communication systems, the primary objective has been to ensure reliable transmission of data, focusing on delivering bit sequences across noisy channels without considering the meaning, context, or purpose of the data being transmitted. Shannon’s mathematical theory of communication focuses on optimizing metrics such as data rate, error rate, and bandwidth efficiency, whilst being agnostic to the ultimate purpose and relevance of the transmitted information [1]. This approach has been widely successful and effective for general communication needs thus far. However, next-generation communication systems, beginning with 6G, require more intelligent and task-aware communication methods to support a wide range of real-time and mission-critical Internet of Things (IoT) applications [2, 3], such as computer vision [4], autonomous driving [5], extended reality (XR) [6], and generative artificial intelligence (AI) [7].
As we move towards these advanced systems, there is a growing recognition that communication should not merely be about transmitting raw data, but about understanding the underlying meaning and purpose of the data. This shift towards task-oriented semantic communication represents a fundamental change in the design of communication networks [8, 9]. Instead of focusing solely on the accurate and efficient transmission of bits, these new approaches aim to ensure that the information most relevant to the specific task or decision-making process is prioritized and delivered with minimal delay and overhead. This is especially critical in 6G-IoT settings, where devices operate under tight bandwidth, latency, and energy constraints, and where privacy and reliability are essential. For example, in a smart city environment [10], rather than transmitting all sensor data from traffic cameras, task-oriented communication focuses on sending only the information necessary to identify and respond to potential hazards or optimize traffic flow in real time.
With the growing success and popularity of deep learning (DL) in various wireless communication applications [11, 12], many emerging task-oriented communication systems have adopted DL approaches to encode task-relevant information to improve task performance and efficiency of the communication system [13, 14, 15]. Nevertheless, most proposed schemes do not focus on quantifying or benchmarking the amount of information that the encoded features retain about the input, primarily due to the computational difficulty of estimating mutual information and the lack of a unified methodology that provides fair and reproducible results. This omission is critical, particularly in IoT and edge computing scenarios, where understanding how much information is kept and whether it is necessary or private has direct implications on trustworthiness, data security, and system interpretability.
@@
To address the aforementioned challenges, we develop a task-oriented communication system based on contrastive learning [17] and disentangled representation learning [18], and we devise a new metric to compute comparative values for a proxy of mutual information between the encoded features and the inputs across different systems, rather than computing the exact mutual information. More specifically, our major contributions are as follows:
  • We derive a lower bound for the mutual information between the encoded features and the target using contrastive learning principles. We show that the contrastive learning based lower bound improves task accuracy and performance compared to traditional cross-entropy based mutual information approximations;
  • We propose a systematic training methodology based on an innovative loss function, designed to extract task-irrelevant information through reconstruction losses while disentangling it from task-relevant features using adversarial methods. This enables the system to prioritize the transmission of task-relevant features while minimizing communication overhead and reducing unnecessary information transmission, thus enhancing privacy;
  • To address the current limitation of lacking a reliable and unified approach for estimating the mutual information between the encoded features and input data, we introduce a new metric named the Information Retention Index (IRI), which serves as a proxy for the mutual information. This metric compares the informativeness and minimality of the encoded features across various task-oriented communication methods, providing deeper insights into system behavior and enabling a more rigorous comparison of their performance;
  • We evaluate our proposed task-oriented communication system in diverse channel conditions. It is tested against several existing task-oriented communication methods, demonstrating improved task performance, enhanced privacy awareness, and reduced amount of irrelevant information across a wide range of transmission scenarios.
@@ II Related Work
DL-based communication systems have shown success in recent years. DeepJSCC (Deep Joint Source-Channel Coding) is a recent advancement in the field of wireless communication that utilizes deep learning to jointly optimize source and channel coding, which are traditionally treated as separate tasks [14]. Unlike conventional methods that rely on separate compression and error-correction codes, DeepJSCC uses neural networks to directly map source data to channel symbols, allowing for an end-to-end optimization of the communication system. DeepJSCC can be trained on a classification task by minimizing the cross-entropy loss, ensuring task-specific performance; however, it does not inherently ensure that only task-relevant features are transmitted.
@@
Wang et al. [20] formulate a privacy-utility trade-off to develop an IB-based privacy-preserving task-oriented communication system against model inversion attacks [21]. This is achieved by striking a balance between the traditional IB-based loss functions similar to the work discussed above and a mean squared error (MSE) based term that aims at maximizing reconstruction distortion. Their results demonstrate improved privacy with minimal impact to task performance.
@@ III-A Notations
Throughout this paper, we use the following notational conventions. Random variables are denoted by uppercase letters, such as $X$, $Y$, and $Z$. Their corresponding realizations (i.e., specific instances) are denoted by lowercase bold letters, such as $\boldsymbol{x}$, $\boldsymbol{y}$, and $\boldsymbol{z}$. The space from which these random variables are drawn is represented by calligraphic letters, such as $\mathcal{X}$, $\mathcal{Y}$ and $\mathcal{Z}$. We denote entropy of a random variable $X$ by $H(X)$. The mutual information between two random variables $X$ and $Y$, is denoted by $I(X;Y)$. We use $I(Z;X|Y)$ to denote the conditional mutual information between $Z$ and $X$ given $Y$. We use the expectation notation $\mathbb{E}[\cdot]$, which refers to the average value of a random variable over a distribution. Most of the symbols in the article are listed in Table I.
@@ III-B System Model
We consider a semantic communication system designed for next-generation 6G-enabled Internet of Things (6G-IoT) networks, where distributed edge devices must transmit task-relevant information to centralized or cloud-based servers under strict constraints on bandwidth, latency, and privacy. In such settings, communication should prioritize the efficient delivery of minimal yet informative representations, discarding irrelevant or sensitive content that is unnecessary for the downstream task.
The transmitter includes a feature extractor and a joint source-channel coding (JSCC) encoder. We collectively refer to these components as the task-relevant encoder. The task-relevant encoder encodes an input image $\boldsymbol{x}\in\mathcal{X}$ into a lower-dimensional feature vector $\boldsymbol{z}\in\mathcal{Z}$. Encoded vector $\boldsymbol{z}$ is then transmitted to a receiver over a noisy wireless channel. The primary objective is to transmit a minimal and informative representation, by discarding task-irrelevant information while ensuring that $\boldsymbol{z}$ contains only the essential information for accurate downstream classification at the receiver.
@@
At the transmitter, the task-relevant encoder encodes input image $\boldsymbol{x}\in\mathbb{R}^{N}$, where $N$ represents the number of pixels in the image (Height $\times$ Width $\times$ Color Channels). The encoder maps this input into a lower-dimensional feature vector $\boldsymbol{z}\in\mathbb{R}^{d}$, where $d$ is the dimension of the encoded feature vector. The encoding function, denoted by $f_{\theta}:\mathbb{R}^{N}\to\mathbb{R}^{d}$, is parameterized by $\theta$, and the encoding process can be expressed as
@@
Feature vector $\boldsymbol{z}$ is then prepared for transmission over the wireless channel by being mapped to channel input symbols. The role of the task-relevant encoder is twofold: encoding the input data into feature representations and preparing them as channel symbols suitable for transmission. The encoded feature vector $\boldsymbol{z}\in\mathbb{R}^{d}$ is transmitted over a wireless channel, which is modeled as an additive white Gaussian noise (AWGN) channel. The channel introduces noise and distortion, and the received signal $\hat{\boldsymbol{z}}\in\mathbb{R}^{d}$ at the receiver is expressed as
@@
At the receiver, typically a cloud or edge server, a classifier, denoted by $q_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{M}$, where $M$ is the number of labels, is parameterized by $\phi$. The classifier maps the received noisy signal $\hat{\boldsymbol{z}}$ to predicted label $\hat{y}\in\mathbb{R}^{M}$ as
@@
TABLE I: Description of Symbols
Symbol Description
$\boldsymbol{y},Y$ Target variable and its realization
$\boldsymbol{x},X$ Input variable and its realization
$\boldsymbol{z},Z$ Encoded feature and its realization
$\hat{\mathbf{z}},\hat{Z}$ Received (noisy) encoded feature and its realization
$\boldsymbol{z}_{1},Z_{1}$ Task-relevant feature and its realization
$\boldsymbol{z}_{2},Z_{2}$ Task-irrelevant feature and its realization
$\hat{\boldsymbol{y}},\hat{Y}$ Predicted label and its realization
$\mathcal{X},\mathcal{Y},\mathcal{Z}$ Input, label, and feature spaces
$H(\cdot)$ Entropy function
$I(\cdot;\cdot),I(\cdot;\cdot\mid\cdot)$ Mutual and conditional mutual information
$\mathbb{E}[\cdot]$ Expectation operator
$\mathbf{I}$ Identity matrix
$p(\cdot)$ Probability distribution
$\mathcal{N}$ Statistical Gaussian distribution
$\mathbf{n}$ Additive Gaussian noise
$f_{\theta}(\cdot)$ Task-relevant encoder, parameterized by $\theta$
$g_{\eta}(\cdot)$ Task-irrelevant encoder, parameterized by $\eta$
$r_{\omega}(\cdot)$ Reconstructor, parameterized by $\omega$
$h_{\psi}(\cdot)$ Projection head , parameterized by $\psi$
$q_{\phi}(\cdot)$ Classifier, parameterized by $\phi$
$D_{\nu}(\cdot)$ Discriminator, parameterized by $\nu$
@@ IV-D Limitations of Variational Information Bottleneck (VIB)
  • Conflicting objectives: The VIB objective which maximizes $I(\hat{Z};X)$ and minimizes $I(\hat{Z};Y)$ leads to a conflicting objective as shown by [32]. If we consider the information diagram [34] presented in Fig. 1, it is evident that the region shaded in blue, namely $I(X;Y;\hat{Z})$ is a subset of $I(\hat{Z};X)$ and $I(\hat{Z};Y)$ and is therefore maximized and minimized simultaneously.
@@ V Proposed Method: CLAD
The core idea of CLAD is to create an end-to-end communication system that explicitly learns to disentangle task-relevant information from task-irrelevant content, enabling both high task accuracy and improved privacy. CLAD integrates supervised contrastive learning, adversarial training, and reconstruction-based supervision into a unified, structured framework that aligns with the MNI principle.
@@
To ensure effective optimization, CLAD adopts a three-stage training strategy, where each stage isolates and refines a different component of the system. This design is intentional as it prevents interference between competing objectives, stabilizes training, and encourages modular reuse of the encoders across stages.
CLAD utilizes two key encoders: the task-relevant encoder, which maps the input into the task-relevant channel codeword $Z_{1}$, and the task-irrelevant encoder , which maps the input into task-irrelevant channel codeword $Z_{2}$. The task performance is optimized through contrastive learning, which aims to maximize the mutual information $I(\hat{Z_{1}};Y)$, ensuring that $\hat{Z_{1}}$ captures the most informative features for downstream classification. The disentanglement is achieved through reconstruction learning to capture task-irrelevant information in $\hat{Z_{2}}$, and adversarial training is utilized to minimize the mutual information $I(\hat{Z_{1}};\hat{Z_{2})}$, thus promoting independence between the two feature representations. The different components and training stages of CLAD are visualized in Fig. 2. These components are optimized together to ensure both high task accuracy and effective disentanglement of information, corresponding to the following maximization objective:
@@ V-A Contrastive Loss for Task-Relevant Features
where $S_{ij}$ represents the similarity score between the projected representations $\boldsymbol{h}_{i}$ and $\boldsymbol{h}_{j}$, $\tau$ is a temperature scaling factor, and $\mathds{1}_{y_{i}=y_{j}}$ is an indicator function that equals 1 if $y_{i}=y_{j}$ (positive pairs) and 0 otherwise (negative pairs). This loss encourages encoder $f_{\theta}(\cdot)$ to learn class-discriminative features, ensuring that the latent representation $\boldsymbol{z}$ captures the necessary information for the downstream classification task. Supervised contrastive learning is presented visually in Fig 3.
@@ V-F Information Retention Index Across Different Methods
Figure 4: Comparison of IRI and MINE estimates across different values of $\beta$ in the VIB objective. Both metrics exhibit similar trends in retained information, but MINE shows significantly higher variance. Shaded regions represent the standard deviation over multiple runs.
Although Mutual Information Neural Estimation (MINE) is a widely-used method for estimating the mutual information [37], it suffers from several well-documented drawbacks that limit its effectiveness for benchmarking in semantic communication systems [38]. MINE is prone to high variance, requires careful tuning of many different hyperparameters, and scales poorly in high-dimensional latent spaces all of which can lead to unstable or misleading estimates. These issues are especially problematic when comparing different models or compression levels under consistent settings. In contrast, IRI allows for consistent, fair benchmarking with minimal assumptions and training overhead. We view IRI as a relative measure, similar to a diagnostic tool, that enables empirical analysis of minimality and informativeness in task-oriented systems. As illustrated in Figure 4, IRI follows the same general trend as MINE but with substantially reduced variance, providing a more stable and reliable metric for evaluating privacy-relevant information retention in 6G-IoT task-oriented communication systems.
We note that while SSIM is effective in image-based applications, it is inherently limited to image data. For non-visual modalities such as text or tabular data, alternative reconstruction-based metrics (e.g., BLEU score) can be adopted to extend the IRI framework to broader task domains.